### v1

@@ 5.1 Proof of the Courcelle’s Theorem for Cliquewidth Using Reduced Evaluation Trees
Paths $P_{1}^{\prime}$ and $P_{2}^{\prime}$ are compatible by 5.15. It remains to show that their combination is equivalent to $P_{1}\otimes P_{2}$. We will prove by induction that for each $i$, the prefixes of $P_{1}\otimes P_{2}$ and $P_{1}^{\prime}\otimes P_{2}^{\prime}$ of length $i$ are equivalent.
  • If $i=0$, every two paths from the root of length 00 are equivalent.
  • If $i>0$, assume that the prefixes of $P_{1}\otimes P_{2}$ and $P_{1}^{\prime}\otimes P_{2}^{\prime}$ of length $i-1$ are equivalent. Let $G_{i}^{\prime},G_{i}$ denote the induced colored subgraph of the prefixes of length $i$ of $P_{1}\otimes P_{2}$ and $P_{1}^{\prime}\otimes P_{2}^{\prime}$, respectively. We will consider two cases: when $z_{i}$ is a set variable and when $z_{i}$ is an individual variable.
    • If $z_{i}$ is a set variable, observe that $G_{i}=G_{i-1}$ and $G_{i}^{\prime}=G_{i-1}^{\prime}$. By the induction assumption $G_{i-1}=G_{i-1}^{\prime}$, so $G_{i}=G_{i}^{\prime}$. Let $S_{1},S_{1}^{\prime}\subseteq V(G_{t_{1}})$ be the labels of the $i$-th node of $P_{1},P_{1}^{\prime}$, respectively, $S_{2},S_{2}^{\prime}\subseteq V(G_{t_{2}})$ be the labels of the $i$-th node of $P_{2},P_{2}^{\prime}$, respectively. Then the label of the $i$-th node of $P_{1}\otimes P_{2}$ is $S_{1}\cup S_{2}$ and the label of the $i$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $S_{1}^{\prime}\cup S_{2}^{\prime}$. For each individual variable $z_{j}$ such that $j<i$, if the label of the $j$-th node of $P_{1}\otimes P_{2}$ is $v\in V(G_{t_{1}})$, then the label of the $j$-th node of $P_{1}$ is $v$ and the label of the $j$-th node of $P_{2}$ is $ext$, since $P_{1}$ and $P_{2}$ were compatible. Then, by 5.15 the label of the $j$-th node of $P_{1}^{\prime}$ is $v^{\prime}\in V(G_{t_{1}})$ and the label of the $j$-th node of $P_{2}^{\prime}$ is $ext$, so the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $v^{\prime}$. Since $P_{1}$ and $P_{1}^{\prime}$ are equivalent, $v\in S_{1}$ if and only if $v^{\prime}\in S_{1}^{\prime}$ and because $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $v\in S_{1}\cup S_{2}$ if and only if $v\in S_{1}$ and $v^{\prime}\in S_{1}^{\prime}\cup S_{2}^{\prime}$ if and only if $v^{\prime}\in S_{1}^{\prime}$, so $v\in S_{1}\cup S_{2}$ if and only if $v^{\prime}\in S_{1}^{\prime}\cup S_{2}^{\prime}$. Symmetrically, if the label of the $j$-th node of $P_{1}\otimes P_{2}$ is $u\in V(G_{t_{2}})$, by the same argument the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $u^{\prime}\in V(G_{t_{2}})$ and $u\in S_{1}\cup S_{2}$ if and only if $u^{\prime}\in S_{1}^{\prime}\cup S_{2}^{\prime}$.
    • If $z_{i}$ is an individual variable, let $v$ be the label of the $i$-th node of $P_{1}\otimes P_{2}$. Without loss of generality, assume that $v\in V(G_{t_{1}})$. Then the label of the $i$-th node of $P_{1}$ is $v$ and the label of the $i$-th node of $P_{2}$ is $ext$, so by 5.15 the label of the $i$-th node of $P_{1}^{\prime}$ is $v^{\prime}\in V(G_{t_{1}})$ and the label of the $i$-th node of $P_{2}^{\prime}$ is $ext$, so the label of the $i$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $v^{\prime}$.
        For a set variable $z_{j}$ such that $j<i$ let $S$ be the label of the $j$-th node of $P_{1}\otimes P_{2}$ and $S^{\prime}$ be the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$. Then $S_{1}\coloneq S\cap V(G_{t_{1}})$ is the label of the $j$-th node of $P_{1}$ and $S_{2}\coloneq S\cap V(G_{t_{2}})$ is the label of the $j$-th node of $P_{2}$. Similarly $S_{1}^{\prime}\coloneq S^{\prime}\cap V(G_{t_{1}})$ is the label of the $j$-th node of $P_{1}^{\prime}$ and $S_{2}^{\prime}\coloneq S^{\prime}\cap V(G_{t_{2}})$ is the label of the $j$-th node of $P_{2}^{\prime}$. Observe that because $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $v\in S$ if and only if $v\in S_{1}$ and similarly $v^{\prime}\in S^{\prime}$ if and only if $v^{\prime}\in S_{1}^{\prime}$. Since $P_{1}$ and $P_{1}^{\prime}$ are equivalent, $v\in S_{1}$ if and only if $v^{\prime}\in S_{1}^{\prime}$. So $v\in S$ if and only if $v^{\prime}\in S^{\prime}$.
        For an individual variable $z_{j}$ such that $j<i$ let $u$ be the label of the $j$-th node of $P_{1}\otimes P_{2}$.
          If $u\in V(G_{t_{2}})$, then $u$ is the label of the $j$-th node of $P_{2}$ and the label of the $j$-th node of $P_{1}$ is $ext$. Then by 5.15 the label of the $j$-th node of $P_{2}^{\prime}$ is $u^{\prime}\in V(G_{t_{2}})$ and the label of the $j$-th node of $P_{1}^{\prime}$ is $ext$, so the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $u^{\prime}$. Then because $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, $u\not=v$ and $u^{\prime}\not=v^{\prime}$, $uv\not\in E(G_{i})$ and $u^{\prime}v^{\prime}\not\in E(G_{i}^{\prime})$.
          If $u\in V(G_{t_{1}})$, then $u$ is the label of the $j$-th node of $P_{1}$ and the label of the $j$-th node of $P_{2}$ is $ext$. Then by 5.15 the label of the $j$-th node of $P_{1}^{\prime}$ is $u^{\prime}\in V(G_{t_{1}})$ and the label of the $j$-th node of $P_{2}^{\prime}$ is $ext$, so the label of the $j$-th node of $P_{1}^{\prime}\otimes P_{2}^{\prime}$ is $u^{\prime}$. Observe that since $P_{1}$ is equivalent to $P_{1}^{\prime}$, $u=v$ if and only if $u^{\prime}=v^{\prime}$, and $uv\in E(G_{t_{1}})$ if and only if $u^{\prime}v^{\prime}\in E(G_{t_{1}})$, so $uv\in E(G_{i})$ if and only if $u^{\prime}v^{\prime}\in E(G_{i}^{\prime})$.∎
Proof of Lemma 5.12.
Now, to prove the Lemma 5.12 we use induction on the height of the subtrees.
  • If the height is 00, let $P_{1},P_{1}^{\prime},P_{2},P_{2}^{\prime}$ be the paths of $T_{1},T_{1}^{\prime},T_{2},T_{1}^{\prime}$ to the root of $F_{1},F_{1}^{\prime},F_{2},F_{2}^{\prime}$, respectively. $T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}$ if and only if $P_{1}\otimes P_{2}$ is equivalent to $P_{1}^{\prime}\otimes P_{2}^{\prime}$. Since $T_{1}\sim T_{1}^{\prime}$ and $T_{2}\sim T_{2}^{\prime}$, $P_{1}$ is equivalent to $P_{1}^{\prime}$ and $P_{2}$ is equivalent to $P_{2}^{\prime}$. Then, by Lemma 5.19, $P_{1}\otimes P_{2}$ is equivalent to $P_{1}^{\prime}\otimes P_{2}^{\prime}$.
@@
  • If the height is greater than 00, for each child subtree $T_{1}^{i}$ of $T_{1}$ there exist a child subtree ${T_{1}^{i}}^{\prime}$ of $T_{1}^{\prime}$ such that $T_{1}^{i}\sim{T_{1}^{i}}^{\prime}$ and for each child subtree $T_{2}^{j}$ of $T_{2}$ there exist a child subtree ${T_{2}^{j}}^{\prime}$ of $T_{2}^{\prime}$ such that $T_{2}^{j}\sim{T_{2}^{j}}^{\prime}$. Then for each child subtree $T_{1}^{i}\otimes T_{2}^{j}$ of $T_{1}\otimes T_{2}$ by 5.18 we know that the paths of $T_{1}^{i},T_{2}^{j}$ to the root of $F_{1},F_{2}$ respectively were compatible. So by 5.15, the paths of ${T_{1}^{i}}^{\prime},{T_{2}^{j}}^{\prime}$ to the root of $F_{1}^{\prime},F_{2}^{\prime}$ respectively were also compatible. Then $T_{1}^{\prime}\otimes T_{2}^{\prime}$ has a child ${T_{1}^{i}}^{\prime}\otimes{T_{2}^{j}}^{\prime}$ which, by the induction assumption, is isomorphic to $T_{1}^{i}\otimes T_{2}^{j}$. Symmetrically for each child subtree ${T_{1}^{i}}^{\prime}$ of $T_{1}^{\prime}$ there exist a child subtree $T_{1}^{i}$ of $T_{1}$ such that $T_{1}^{i}\sim{T_{1}^{i}}^{\prime}$ and for each child subtree ${T_{2}^{j}}^{\prime}$ of $T_{2}$ there exist a child subtree ${T_{2}^{j}}^{\prime}$ of $T_{2}^{\prime}$ such that $T_{2}^{j}\sim{T_{2}^{j}}^{\prime}$. Then, by the same argument, for each child subtree ${T_{1}^{i}}^{\prime}\otimes{T_{2}^{j}}^{\prime}$ of $T_{1}^{\prime}\otimes T_{2}^{\prime}$ there exist a child subtree $T_{1}^{i}\otimes T_{2}^{j}$ of $T_{1}\otimes T_{2}$ which is isomorphic to ${T_{1}^{i}}^{\prime}\otimes{T_{2}^{j}}^{\prime}$. Therefore, $T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}$.∎
A direct application of Lemma 5.12 is the following corollary.
Corollary 5.20.
@@
In other words, we proved that the product of two reduced partial evaluation trees is isomorphic to the product of unreduced ones. Having this, we are ready to prove the main theorem.
Proof of Theorem 5.1.
@@
For each $t\in V(\mathcal{D})$ we can construct a reduced partial evaluation tree $F_{t}^{-}$ for $\Phi$ which is isomorphic to $F_{t}$ by induction as follows:
@@
    Note that, by 5.10,
    $|F_{t}^{\prime}|\leq|F_{t_{1}}^{-}|\cdot|F_{t_{2}}^{-}|\leq\left(\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$}\right)^{2}=\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1.$}$
@@
    If two configurations of leaves of $F_{t^{\prime}}$ were isomorphic, they stay isomorphic after recoloring. However, some non-isomorphic configurations may become isomorphic, so we have to reduce them taking into account the new coloring $\mu$. Then
@@
    If two configurations of leaves of $F_{t^{\prime}}$ were isomorphic, that means that the induced subgraphs of these configurations were isomorphic, including the coloring of vertices, so after adding the new edges they stay isomorphic. However, some non-isomorphic configurations may become isomorphic, so we have to reduce them taking into account the new edges. Then
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1.$}\qed$
5.2 Generalization to Diverse Vertex Problems
Let $\mathcal{P}$ be a vertex problem expressible in $MSO_{1}$ logic and $\Phi=\exists_{z_{1}}\Phi^{\prime}(z_{1})$ be an $MSO_{1}$ formula for this problem, where the value of $z_{1}$ represents the set $S$ such that $(G,S)\in\mathcal{P}$. Assume that $\Phi^{\prime}$ is in the prenex normal form, so $\Phi^{\prime}=\exists_{z_{1}\subseteq V(G)}\ldots\Psi(z_{1},\ldots,z_{q})$, where $\Psi(z_{1},\ldots,z_{q})$ is the quantifier-free part of $\Phi$.
@@ 5.2 Generalization to Diverse Vertex Problems
Let $\mathcal{P}$ be a vertex problem expressible in $MSO_{1}$ logic. There exists a monotone dynamic programming core $\mathfrak{C}$ that (together with a proper $\mathfrak{C}$-vertex-membership function $\rho$) solves $\mathcal{P}$ s.t.
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}\tau(\mathfrak{C},\mathcal{D})=|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$,}$
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}Size(\mathfrak{C},\mathcal{D})=2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
@@
A reduced partial evaluation tree $F_{t}^{-}$ for $\Phi$ and $t\in V(\mathcal{D})$ can be viewed as a set $\mathcal{F}_{t}^{-}$ of non-isomorphic trees of height $q$ (subtrees of nodes of level 1 of $F_{t}^{-}$) with roots labeled by the choices of the value of $z_{1}$ and
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|\mathcal{F}_{t}^{-}|\leq 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f_{1}^{-},f_{2}^{-})|f^{-}\in\mathcal{F}_{t}^{-},f_{1}^{-}\in\mathcal{F}_{t_{1}}^{-},f_{2}^{-}\in\mathcal{F}_{t_{2}}^{-},f^{-}=reduce(f_{1}^{-}\otimes f_{2}^{-}))\},$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f^{{}^{\prime}-})|f^{-}\in\mathcal{F}_{t}^{-},f^{{}^{\prime}-}\in\mathcal{F}_{t^{\prime}}^{-},f^{-}=reduce(f^{{}^{\prime}-})\}.$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f^{{}^{\prime}-})|f^{-}\in\mathcal{F}_{t}^{-},f^{{}^{\prime}-}\in\mathcal{F}_{t^{\prime}}^{-},f^{-}=reduce(f^{{}^{\prime}-})\}.$
Observe that for each $t\in V(\mathcal{D})$,
$|Process_{\mathfrak{C},\mathcal{D}}(t)|\leq\prod_{i=1}^{\delta(t)}|\mathcal{F}_{t_{i}}^{-}|\leq{\left(\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}\right)}^{2}=\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$.}$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q+1$}.$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|F_{t}^{-}|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}\leq\left(\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}\right)^{2}=\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log k))}}}}}\end{array}\right\}\text{\scriptsize$q$}.$
It remains to prove that $\mathfrak{C}$ with a proper membership function $\rho$ solves $\mathcal{P}$. For $f^{-}\in\mathcal{F}_{t}^{-}$ let
$\rho(f^{-})=\begin{cases}0&\text{if the root of }f^{-}\text{ is labeled by }\emptyset,\\ 1&\text{otherwise}.\end{cases}$

### v2

@@ 5.1 Proof of Courcelle’s Theorem for Cliquewidth Using Reduced Evaluation Trees
Assume the root configurations of $T_{1}$ and $T_{2}$ are compatible. By ˜5.14, the root configurations of $T_{1}$ and $T_{1}^{\prime}$ are the same, similarly the root configurations of $T_{2}$ and $T_{2}^{\prime}$, so the root configurations of $T_{1}^{\prime}$ and $T_{2}^{\prime}$ are also compatible. Then, by ˜5.16 $T_{1}\otimes T_{2}$ is a subtree of $F_{1}\otimes F_{2}$ and $T_{1}^{\prime}\otimes T_{2}^{\prime}$ is a subtree of $F_{1}^{\prime}\otimes F_{2}^{\prime}$.
@@
It remains to prove that $T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}$}. We do it by induction on the height of the subtrees. As $G_{t_{1}}$ and $G_{t_{2}}$ are disjoint, by ˜5.16 the root configurations of $T_{1}\otimes T_{2}$ and $T_{1}^{\prime}\otimes T_{2}^{\prime}$ are the same and equal to the union of the root configuration of $T_{1}$ and $T_{2}$. For subtrees of height 0, that means they are isomorphic. Assume now that the height is greater than 0. It suffices to prove that for each child subtree of $T_{1}\otimes T_{2}$ in $F_{1}\otimes F_{2}$, there exists an isomorphic child subtree of $T_{1}^{\prime}\otimes T_{2}^{\prime}$ in $F_{1}^{\prime}\otimes F_{2}^{\prime}$. Proof of the converse implication will be symmetric. For each pair of children subtrees $T_{1}^{i},T_{2}^{j}$ of $T_{1}$ and $T_{2}$, respectively, there exists a pair of children subtrees $T_{1}^{\prime i},T_{2}^{\prime j}$ of $T_{1}^{\prime}$ and $T_{2}^{\prime}$, respectively, such that $T_{1}^{\prime i}\sim T_{1}^{i}$ and $T_{2}^{\prime j}\sim T_{2}^{j}$. If $T_{1}^{i}\otimes T_{2}^{j}$ is a child subtree of $T_{1}\otimes T_{2}$ in $F_{1}\otimes F_{2}$, the configurations of $T_{1}^{i}$ and $T_{2}^{j}$ are compatible. Then, by ˜5.14 the configurations of $T_{1}^{\prime i}$ and $T_{2}^{\prime j}$ are also compatible, so $T_{1}^{\prime i}\otimes T_{2}^{\prime j}$ is a child subtree of $T_{1}^{\prime}\otimes T_{2}^{\prime}$ in $F_{1}^{\prime}\otimes F_{2}^{\prime}$. By the induction assumption $T_{1}^{\prime i}\otimes T_{2}^{\prime j}\sim T_{1}^{i}\otimes T_{2}^{j}$ as subtrees of $F_{1}\otimes F_{2}$ and $F_{1}^{\prime}\otimes F_{2}^{\prime}$ of height one smaller than $T_{1}\otimes T_{2},T_{1}^{\prime}\otimes T_{2}^{\prime}$. Therefore, $T_{1}\otimes T_{2}\sim T_{1}^{\prime}\otimes T_{2}^{\prime}$}. ∎
@@
In other words, Lemma˜5.12 says the product of two reduced partial evaluation trees is isomorphic to the product of unreduced ones. Having this, we are ready to prove the main theorem.
Proof of Theorem˜5.1.
@@
For each $t\in V(\mathcal{D})$, we can construct a reduced partial evaluation tree $F_{t}^{-}$ for $\Phi$ which is isomorphic to $F_{t}$ by induction as follows:
@@
    Note that, by ˜5.10,
    $|F_{t}^{\prime}|\leq|F_{t_{1}}^{-}|\cdot|F_{t_{2}}^{-}|\leq\left(\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q+1$}\right)^{2}=\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q+1.$}$
@@
    If two configurations of leaves of $F^{-}_{t^{\prime}}$ were isomorphic, they stay isomorphic after recoloring. However, some non-isomorphic configurations may become isomorphic, so we have to reduce them by taking into account the new coloring $\mu$. Then
@@
    If two configurations of leaves of $F^{-}_{t^{\prime}}$ were isomorphic, that means that the induced subgraphs of these configurations were isomorphic, including the coloring of vertices, so after adding the new edges, they stay isomorphic. However, some non-isomorphic configurations may become isomorphic, so we have to reduce them, taking into account the new edges. Then
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q+1.$}$
The correctness of the above approach is implied by Lemma˜5.12. ∎
5.2 Dynamic Programming Core for MSO1-expressible Vertex Problems
@@
Let $\mathcal{P}$ be a vertex problem expressible in $MSO_{1}$ logic and $\Phi$ be the formula for this problem with $q$ quantified variables; $q_{v}$ of them being individual variables. There exists a monotone dynamic programming core $\mathfrak{C}$ that (together with a proper $\mathfrak{C}$-vertex-membership function $\rho$) solves $\mathcal{P}$ such that
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}\tau(\mathfrak{C},\mathcal{D})=|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q+1$,}$
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}Size(\mathfrak{C},\mathcal{D})=2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q$,}$
where $\omega$ is the width of $\mathcal{D}$.
@@
Let $\mathcal{P}$ be a vertex problem expressible in $MSO_{1}$ logic and $\Phi=\exists_{z_{1}}\Phi^{\prime}(z_{1})$ be an $MSO_{1}$ formula for this problem in the prenex normal form, where the value of $z_{1}$ represents the set $S$ such that $(G,S)\in\mathcal{P}$. A reduced partial evaluation tree $F_{t}^{-}$ for $\Phi$ and $t\in V(\mathcal{D})$ can be viewed as a set $\mathcal{F}_{t}^{-}$ of non-isomorphic trees of height q1q-1 (subtrees of nodes of level 1 of $F_{t}^{-}$) with roots labeled by the choices of the value of $z_{1}$ and
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|\mathcal{F}_{t}^{-}|\leq 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f_{1}^{-},f_{2}^{-})\in\mathcal{F}_{t}^{-}\times\mathcal{F}_{t_{1}}^{-}\times\mathcal{F}_{t_{2}}^{-}:f^{-}=reduce(f_{1}^{-}\otimes f_{2}^{-}))\},$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f^{{}^{\prime}-})\in\mathcal{F}_{t}^{-}\times\mathcal{F}_{t^{\prime}}^{-}:f^{-}=reduce(f^{{}^{\prime}-})\}.$
@@
    $Process_{\mathfrak{C},\mathcal{D}}(t)=\{(f^{-},f^{{}^{\prime}-})\in\mathcal{F}_{t}^{-}\times\mathcal{F}_{t^{\prime}}^{-}:f^{-}=reduce(f^{{}^{\prime}-})\}.$
Observe that for each $t$,
$|Process_{\mathfrak{C},\mathcal{D}}(t)|\leq\prod_{i=1}^{\delta(t)}|\mathcal{F}_{t_{i}}^{-}|\leq\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q+1$.}$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|V(\mathcal{D})|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q+1$}.$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q$.}$
@@
$\left.\kern-1.2pt\begin{array}[]{@{}c@{}}|\mathcal{F}_{t}^{-}|\cdot 2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q$}\leq\left.\kern-1.2pt\begin{array}[]{@{}c@{}}2^{\cdot^{\cdot^{\cdot^{\scriptstyle 2^{\mathcal{O}(q_{v}(q+\log\omega))}}}}}\end{array}\right\}\text{\scriptsize$q$}.$
It remains to define a proper membership function. For $f^{-}\in\mathcal{F}_{t}^{-}$ let
[…]
Claim 5.21.
Dynamic programming core $\mathfrak{C}$ together with the membership function $\rho$ defined as above solves $\mathcal{P}$.
Proof of ˜5.21.